OpenStudy (anonymous):

I will give you a medal and be your fan How do you convert an improper fraction into a mixed number?

OpenStudy (anonymous):

Writing Improper Fractions as Mixed Numbers To write an improper fraction as a mixed number, divide the numerator (top part) by the denominator (bottom part). The quotient is the whole number, and the remainder is the numerator. How would you express 17/4 as a mixed number? Divide the numerator by the denominator: 17 ÷ 4 = 4, with a remainder of 1 The quotient, 4, is the whole number. The remainder, 1, is the numerator. The denominator remains 4. 17/4 = 41/4

OpenStudy (anonymous):

Thank you,mabey know that I know this information I can hopefully fully understand improper fractions and mixed numbers

OpenStudy (anonymous):

Welcome , my pleasure